OpenStudy (anonymous):

you play a game in which two dice are rolled if a sum of 7appears, you win 10 otherwise you lose $ 2.00. if you intend to play this game for a long time should you expect to play this game for a lone time should you expeet to make money. lose money or come out about even?

OpenStudy (anonymous):

take the amount you win times the probability you win it and add it up

OpenStudy (anonymous):

probability you roll a seven is \(\frac{1}{6}\) and so the probabablity you do not roll a seven is evidently \(1-\frac{1}{6}+\frac{5}{6}\) so you multiply and add \[10\times \frac{1}{6}-2\times \frac{5}{6}\] and see what you get for your expected value

OpenStudy (anonymous):

another damned typo i means \(1-\frac{1}{6}=\frac{5}{6}\)

OpenStudy (unklerhaukus):

There are 36 possible outcomes there are 6 favorable outcomes \[\langle x \rangle = \sum x P(x)\] \[=\frac{1}{N} \sum xN(x)\] \[=\frac{1}{36} ($10\times 6 -$2\times30)\] \[=\frac{$0}{36}\] \[\langle x \rangle =$0\] you can expect to break even over a long period